# Keep It on a Leash: Controllable Pseudo-label Generation Towards Realistic Long-Tailed Semi-Supervised Learning

### v5

@@ Appendix B Proof of Theorem 1
where $\circ$ denotes the function composition operator, $\mathbb{E}_{(x,y,\mu)}$ denotes the expectation over $x$, $y$, and $\mu$, $\mu$ denotes the Rademacher variable, $\sup_{f\in\mathcal{F}}$ denotes the supremum (or least upper bound) over the function set $\mathcal{F}$ of model $f$. The second line holds because of the Rademacher vector contraction inequality [43].
Then, we proceed with the proof by showing that the one direction $\sup_{f\in\mathcal{F}}R(f)-\widehat{R}(f)$ is bounded with probability at least $1-\upsilon/2$, and the other direction $\sup_{f\in\mathcal{F}}\widehat{R}(f)-R(f)$ can be proved similarly. Note that replacing a sample $(x_{i},y_{i})$ leads to a change of $\sup_{f\in\mathcal{F}}R(f)-\widehat{R}(f)$ at most $\frac{U}{N+M}$ due to the fact that $\ell_{la}$ is bounded by $U$. According to the McDiarmid’s inequality [38], for any $\upsilon>0$, with probability at least $1-\upsilon/2$, we have:
@@
According to the result in  [38] that shows $\mathbb{E}\left[\sup_{f\in\mathcal{F}}R(f)-\widehat{R}(f)\right]\leq 2\mathcal{R}_{N+M}(\mathcal{F})$, and further considering the other direction $\sup_{f\in\mathcal{F}}\widehat{R}(f)-R(f)$, with probability at least $1-\upsilon$, we have:
@@ Appendix C Dataset Details
We perform our experiments on four widely-used datasets (CIFAR-10-LT [39], CIFAR-100-LT [39], Food-101-LT [40], and SVHN-LT [41]), following the main experimental settings in FreeMatch [8] and SimPro [13], details are as below.
@@ Appendix D Implementation Details
We follow the default settings and hyperparameters in USB, i.e., the batch size of labeled data $B_{l}$ is set to 64, while unlabeled data $B_{u}$ is set to 7 times $B_{l}$, and the confidence threshold $\tau$ is set to 0.95. Moreover, we use the WRN-28-2 [44] architecture, the SGD optimizer with momentum 0.9, and weight decay 5e-4 for training. We use a cosine learning rate decay [45] scheme, which sets the learning rate to $\eta\cos\left(\frac{7\pi t}{16T}\right)$, where the initial learning rate $\eta$ is set to 0.03, $t$ is the current training step, and the total number of training steps $T$ is set to $2^{18}$. We repeat each experiment with three different random seeds (i.e., 0, 1, and 2) and report the mean and standard deviation of the performances. We conduct the experiments on a single NVIDIA RTX 4090 GPU using PyTorch v2.3.1.
@@ E.1 Comparison of Macro-F1 on CIFAR-10-LT and CIFAR-100-LT
[…]
@@ E.2 Evaluation on Audio Modality
FreeMatch [8] SimPro [13] Ours 68.13 68.75 69.13
While our method is evaluated on vision tasks, the core principles of CPG can be extended to other modalities. Here, we further evaluate our method on the audio modality. Specifically, for audio data, this adaptation requires two key modifications: (i) replacing vision-specific backbones (e.g., WRN-28-2 [44]) with audio-compatible models (e.g., HuBERT [46]), and (ii) substituting image augmentations with audio-specific ones. To demonstrate this adaptability, we perform our CPG on ESC-50-LT [47], a standard benchmark for environmental sound classification. As demonstrated in Table 8, our method achieves state-of-the-art performance on this audio benchmark, confirming its effectiveness beyond the visual modality. These results strongly support the generalizability of CPG’s framework across different modalities.
@@ E.3 Evaluation under Noisy Scenarios
Noise rate 0% 10% 20% 30% $\Delta$ (0% $\rightarrow$ 30%) FreetMatch [8] 66.33 60.33 56.24 51.29 -15.04 SimPro [13] 64.73 59.86 44.54 34.80 -29.93 Ours 82.33 75.95 71.82 69.87 -12.46
To comprehensively evaluate our method’s robustness under noisy scenarios, we conduct additional experiments on CIFAR-10-LT with $(N_{max}=400,M_{max}=4600,\gamma=100)$, under varying noise rates (i.e., 0%, 10%, 20%, and 30%). The labeled data follows a long-tailed distribution, while the unlabeled data adheres to an arbitrary distribution. As demonstrated in Table 9, our method consistently outperforms baselines in noisy settings, achieving the highest accuracy across all noise rates with minimal performance degradation. Notably, when noise rate increases from 0% to 30%, our method shows a degradation ($\Delta$) of only 12.46 pp, compared to 15.04 pp for FreeMatch [8] and 29.93 pp for SimPro [13]. This demonstrates our method’s strong robustness to label noise.
@@ E.4 Evaluation on Large-scale and Realistic Datasets
[…]
@@ E.5 Evaluation under Extreme Settings
Method Setting A1 Setting A2 Setting B FreetMatch [8] 63.86 45.75 55.92 SimPro [13] 31.01 15.20 50.98 Ours 69.56 49.28 77.78
@@ E.6 Evaluation using Other Architectures
Method Arbitrary Inverse FreetMatch [8] 47.68 48.28 SimPro [13] 43.67 51.26 Ours 55.53 58.49
@@ F.1 Ablation Study on the Confidence Threshold
Method Arbitrary Inverse Consistent FreetMatch [8] 66.33 68.10 69.42 SimPro [13] 64.73 65.25 64.37 Ours ($\tau=0.75$) 81.35 78.26 76.20 Ours ($\tau=0.85$) 81.89 80.85 77.79 Ours ($\tau=0.95$) 82.33 82.32 78.35
@@ Appendix G Relation to Existing Methods
Here, we summarize the difference between our CPG and existing methods (i.e., FixMatch [7], FlexMatch [29], UPS [48], CADR [49], FlexDA [50], DASO [51]).
@@
While CPG adopts confidence-based filtering similar to FixMatch [7] and FlexMatch [29], its key innovation is the controllable self-reinforcing optimization cycle, which incorporates reliable pseudo-labels from the unlabeled dataset into the labeled dataset rather than employing consistency regularization, distinguishing it from existing methods. This cycle operates in three steps: (i) expanding the labeled dataset with reliable pseudo-labels, (ii) constructing a Bayes-optimal classifier, and (iii) iteratively improving pseudo-label quality, forming a theoretically grounded feedback loop. Additionally, CPG ensures full sample utilization via an auxiliary branch and reduces label noise accumulation through a voting-based stabilization technique, which enforces consistent pseudo-label assignments across training steps. As evidenced by the ablation studies in Table 5, these components individually contribute significant performance improvements, with the auxiliary branch and optimization cycle yielding average gains of 1.82 pp and 6.97 pp, respectively.
@@
The goals of UPS [48] and CADR [49] differ from ours. UPS targets poor model calibration in SSL without distribution mismatch and CADR addresses distribution mismatch due to label missing not at random, while our work focuses on handling the unknown and arbitrary unlabeled data distribution. Specifically, UPS mitigates high-confidence incorrect pseudo-labels caused by poor model calibration through uncertainty estimation techniques like MC-Dropout [52], and CADR focuses on resolving classifier bias due to label missing not at random by employing class-aware propensity estimation and dynamic threshold adjustment for pseudo-label selection. Unlike these existing techniques that rely on uncertainty estimation or threshold adjustment, we propose a novel controllable self-reinforcing optimization cycle that operates without distribution estimation or correction, introducing distinct conceptual and methodological innovations. More importantly, our framework is specifically designed to address real-world scenarios with unknown and arbitrary unlabeled data distributions.
@@
While FlexDA [50] relies on estimating the unlabeled data distribution for distribution alignment and pseudo-label generation, our core contribution introduces a controllable self-reinforcing optimization cycle that entirely bypasses unlabeled data distribution estimation. This approach is fundamentally different from FlexDA’s dynamic distribution alignment. Specifically, we iteratively expand the labeled dataset with reliable pseudo-labels, train a Bayes-optimal classifier via logit adjustment on the updated labeled dataset without using the pseudo-label distribution. More importantly, we further theoretically prove that this optimization cycle can significantly reduce the generalization error, whereas FlexDA lacks such theoretical guarantees. Moreover, our auxiliary branch leverages all available samples (without a confidence threshold) to enhance feature learning while preventing the primary branch classifier from error pseudo-labels. Notably, prior methods (e.g., FlexDA [50], SimPro [50], UPS [48], DASO [51]) lack class-aware enhancements for tail classes.

### v6

@@ Appendix B Proof of Theorem 1
where $\circ$ denotes the function composition operator, $\mathbb{E}_{(x,y,\mu)}$ denotes the expectation over $x$, $y$, and $\mu$, $\mu$ denotes the Rademacher variable, $\sup_{f\in\mathcal{F}}$ denotes the supremum (or least upper bound) over the function set $\mathcal{F}$ of model $f$. The second line holds because of the Rademacher vector contraction inequality [virc].
Then, we proceed with the proof by showing that the one direction $\sup_{f\in\mathcal{F}}R(f)-\widehat{R}(f)$ is bounded with probability at least $1-\upsilon/2$, and the other direction $\sup_{f\in\mathcal{F}}\widehat{R}(f)-R(f)$ can be proved similarly. Note that replacing a sample $(x_{i},y_{i})$ leads to a change of $\sup_{f\in\mathcal{F}}R(f)-\widehat{R}(f)$ at most $\frac{U}{N+M}$ due to the fact that $\ell_{la}$ is bounded by $U$. According to the McDiarmid’s inequality [FML], for any $\upsilon>0$, with probability at least $1-\upsilon/2$, we have:
@@
According to the result in  [FML] that shows $\mathbb{E}\left[\sup_{f\in\mathcal{F}}R(f)-\widehat{R}(f)\right]\leq 2\mathcal{R}_{N+M}(\mathcal{F})$, and further considering the other direction $\sup_{f\in\mathcal{F}}\widehat{R}(f)-R(f)$, with probability at least $1-\upsilon$, we have:
@@ Appendix C Dataset Details
We perform our experiments on four widely-used datasets (CIFAR-10-LT [CIFAR], CIFAR-100-LT [CIFAR], Food-101-LT [Food], and SVHN-LT [SVHN]), following the main experimental settings in FreeMatch [FreeMatch] and SimPro [SimPro], details are as below.
@@ Appendix D Implementation Details
We follow the default settings and hyperparameters in USB, i.e., the batch size of labeled data $B_{l}$ is set to 64, while unlabeled data $B_{u}$ is set to 7 times $B_{l}$, and the confidence threshold $\tau$ is set to 0.95. Moreover, we use the WRN-28-2 [WRN] architecture, the SGD optimizer with momentum 0.9, and weight decay 5e-4 for training. We use a cosine learning rate decay [SGDR] scheme, which sets the learning rate to $\eta\cos\left(\frac{7\pi t}{16T}\right)$, where the initial learning rate $\eta$ is set to 0.03, $t$ is the current training step, and the total number of training steps $T$ is set to $2^{18}$. We repeat each experiment with three different random seeds (i.e., 0, 1, and 2) and report the mean and standard deviation of the performances. We conduct the experiments on a single NVIDIA RTX 4090 GPU using PyTorch v2.3.1.
@@ E.1 Comparison of Macro-F1 on CIFAR-10-LT and CIFAR-100-LT
Scenario Method CIFAR-10-LT Avg. CIFAR-100-LT Avg. $\gamma=100$ $\gamma=10$ Arbitrary Inverse Consistent Arbitrary Inverse Consistent SSL FixMatch [FixMatch] 53.52 54.24 66.47 58.07 44.58 44.85 46.46 45.30 FreetMatch [FreeMatch] 66.00 67.00 68.04 67.01 44.13 43.57 43.84 43.85 SoftMatch [SoftMatch] 62.49 66.66 71.53 66.89 45.67 46.73 46.91 46.44 ReaLTSSL $\dagger [ACR$] 59.01 64.03 70.25 64.43 45.69 48.89 41.92 45.50 SimPro [SimPro] 61.22 62.11 57.81 60.38 39.98 41.54 41.87 41.13 CDMAD [CDMAD] 61.73 62.88 62.11 62.24 32.97 35.80 36.68 35.15 Ours 81.99 82.31 78.25 80.85 51.31 51.25 50.41 50.99
@@ E.2 Evaluation on Audio Modality
FreeMatch [FreeMatch] SimPro [SimPro] Ours 68.13 68.75 69.13
While our method is evaluated on vision tasks, the core principles of CPG can be extended to other modalities. Here, we further evaluate our method on the audio modality. Specifically, for audio data, this adaptation requires two key modifications: (i) replacing vision-specific backbones (e.g., WRN-28-2 [WRN]) with audio-compatible models (e.g., HuBERT [HuBERT]), and (ii) substituting image augmentations with audio-specific ones. To demonstrate this adaptability, we perform our CPG on ESC-50-LT [ESC], a standard benchmark for environmental sound classification. As demonstrated in Table 8, our method achieves state-of-the-art performance on this audio benchmark, confirming its effectiveness beyond the visual modality. These results strongly support the generalizability of CPG’s framework across different modalities.
@@ E.3 Evaluation under Noisy Scenarios
Noise rate 0% 10% 20% 30% $\Delta$ (0% $\rightarrow$ 30%) FreetMatch [FreeMatch] 66.33 60.33 56.24 51.29 -15.04 SimPro [SimPro] 64.73 59.86 44.54 34.80 -29.93 Ours 82.33 75.95 71.82 69.87 -12.46
To comprehensively evaluate our method’s robustness under noisy scenarios, we conduct additional experiments on CIFAR-10-LT with $(N_{max}=400,M_{max}=4600,\gamma=100)$, under varying noise rates (i.e., 0%, 10%, 20%, and 30%). The labeled data follows a long-tailed distribution, while the unlabeled data adheres to an arbitrary distribution. As demonstrated in Table 9, our method consistently outperforms baselines in noisy settings, achieving the highest accuracy across all noise rates with minimal performance degradation. Notably, when noise rate increases from 0% to 30%, our method shows a degradation ($\Delta$) of only 12.46 pp, compared to 15.04 pp for FreeMatch [FreeMatch] and 29.93 pp for SimPro [SimPro]. This demonstrates our method’s strong robustness to label noise.
@@ E.4 Evaluation on Large-scale and Realistic Datasets
Scenario Method 32×32 64×64 SSL FixMatch [FixMatch] 29.57 37.40 FreetMatch [FreeMatch] 31.37 39.67 SoftMatch [SoftMatch] 31.50 39.40 ReaLTSSL $\dagger [ACR$] 38.47 47.61 SimPro [SimPro] 43.31 46.93 CDMAD [CDMAD] 15.46 21.92 Ours 44.58 50.43
@@ E.5 Evaluation under Extreme Settings
Method Setting A1 Setting A2 Setting B FreetMatch [FreeMatch] 63.86 45.75 55.92 SimPro [SimPro] 31.01 15.20 50.98 Ours 69.56 49.28 77.78
@@ E.6 Evaluation using Other Architectures
Method Arbitrary Inverse FreetMatch [FreeMatch] 47.68 48.28 SimPro [SimPro] 43.67 51.26 Ours 55.53 58.49
@@ F.1 Ablation Study on the Confidence Threshold
Method Arbitrary Inverse Consistent FreetMatch [FreeMatch] 66.33 68.10 69.42 SimPro [SimPro] 64.73 65.25 64.37 Ours ($\tau=0.75$) 81.35 78.26 76.20 Ours ($\tau=0.85$) 81.89 80.85 77.79 Ours ($\tau=0.95$) 82.33 82.32 78.35
@@ Appendix G Relation to Existing Methods
Here, we summarize the difference between our CPG and existing methods (i.e., FixMatch [FixMatch], FlexMatch [FlexMatch], UPS [UPS], CADR [CADR], FlexDA [FlexDA], DASO [DASO]).
@@
While CPG adopts confidence-based filtering similar to FixMatch [FixMatch] and FlexMatch [FlexMatch], its key innovation is the controllable self-reinforcing optimization cycle, which incorporates reliable pseudo-labels from the unlabeled dataset into the labeled dataset rather than employing consistency regularization, distinguishing it from existing methods. This cycle operates in three steps: (i) expanding the labeled dataset with reliable pseudo-labels, (ii) constructing a Bayes-optimal classifier, and (iii) iteratively improving pseudo-label quality, forming a theoretically grounded feedback loop. Additionally, CPG ensures full sample utilization via an auxiliary branch and reduces label noise accumulation through a voting-based stabilization technique, which enforces consistent pseudo-label assignments across training steps. As evidenced by the ablation studies in Table 5, these components individually contribute significant performance improvements, with the auxiliary branch and optimization cycle yielding average gains of 1.82 pp and 6.97 pp, respectively.
@@
The goals of UPS [UPS] and CADR [CADR] differ from ours. UPS targets poor model calibration in SSL without distribution mismatch and CADR addresses distribution mismatch due to label missing not at random, while our work focuses on handling the unknown and arbitrary unlabeled data distribution. Specifically, UPS mitigates high-confidence incorrect pseudo-labels caused by poor model calibration through uncertainty estimation techniques like MC-Dropout [MC-Dropout], and CADR focuses on resolving classifier bias due to label missing not at random by employing class-aware propensity estimation and dynamic threshold adjustment for pseudo-label selection. Unlike these existing techniques that rely on uncertainty estimation or threshold adjustment, we propose a novel controllable self-reinforcing optimization cycle that operates without distribution estimation or correction, introducing distinct conceptual and methodological innovations. More importantly, our framework is specifically designed to address real-world scenarios with unknown and arbitrary unlabeled data distributions.
@@
While FlexDA [FlexDA] relies on estimating the unlabeled data distribution for distribution alignment and pseudo-label generation, our core contribution introduces a controllable self-reinforcing optimization cycle that entirely bypasses unlabeled data distribution estimation. This approach is fundamentally different from FlexDA’s dynamic distribution alignment. Specifically, we iteratively expand the labeled dataset with reliable pseudo-labels, train a Bayes-optimal classifier via logit adjustment on the updated labeled dataset without using the pseudo-label distribution. More importantly, we further theoretically prove that this optimization cycle can significantly reduce the generalization error, whereas FlexDA lacks such theoretical guarantees. Moreover, our auxiliary branch leverages all available samples (without a confidence threshold) to enhance feature learning while preventing the primary branch classifier from error pseudo-labels. Notably, prior methods (e.g., FlexDA [FlexDA], SimPro [FlexDA], UPS [UPS], DASO [DASO]) lack class-aware enhancements for tail classes.